# Enhancing Fluorescence Lifetime Parameter Estimation Accuracy with Differential Transformer Based Deep Learning Model Incorporating Pixelwise Instrument Response Function

### v1

@@ 4 Discussion and Conclusion
In this study, we introduced MFliNet, a novel deep learning model based on the DIFF Transformer architecture, to address the challenges of accurate FLI parameter estimation, particularly in complex and variable biological environments. Our results, as illustrated in Figure 2, demonstrate shifts in IRFs at varying heights, highlighting how each organ’s unique geometry and composition contribute to IRF offsets. This variation in the IRF offset underscores the challenge of accurately estimating the FLI parameters and the necessity for MFliNet, which can adapt to these complexities. The integration of pixel-wise IRF analysis within MFliNet specifically addresses the effects of surface irregularities on early photon arrival times, which is often overlooked in other DL models. The differential attention mechanism within the DIFF Transformer enhances the model’s ability to focus on relevant features while suppressing noise, leading to improved performance.
@@
Funding

### v2

@@ 4 Discussion and Conclusion
In this study, we introduced MFliNet, a novel deep learning model based on the DIFF Transformer architecture, to address the challenges of accurate FLI parameter estimation, particularly in complex and variable biological environments. Our results, as illustrated in Figure 2, demonstrate shifts in IRFs at varying heights, highlighting how each organ’s unique geometry and composition contribute to IRF offsets. This variation in the IRF offset underscores the challenge of accurately estimating the FLI parameters and the necessity for MFliNet, which can adapt to these complexities. The integration of pixel-wise IRF analysis within MFliNet specifically addresses the effects of surface irregularities on early photon arrival times, which is often overlooked in other DL models. By comparing MFliNet with a standard transformer model of same architecture and parameters, we demonstrated that the differential attention mechanism inherent in the DIFF Transformer enhances the model’s ability to focus on relevant features while suppressing noise, leading to improved FLI parameter estimation.
@@
Funding This work was supported by the National Institutes of 267 Health (NIH) Grants R01CA237267, R01CA250636, R01CA271371, & R01CA250636-02S1.